A sled is pulled 65 meters at an angle of 45 degrees. If 900 J of work was used how much force in N was used to pull the sled?


1
Expert's answer
2020-11-02T09:21:55-0500

By the definition of the work done we have:


"W=Fscos\theta,""F=\dfrac{W}{scos\theta}=\dfrac{900\ J}{65\ m\cdot cos45^{\circ}}=19.6\ N."

Answer:

"F=19.6\ N."